# JANET: Joint Adaptive predictioN-region Estimation for Time-series

### v1

@@ Abstract
Keywords: Joint Prediction Region, Conformal Prediction, Uncertainty in Time Series Forecasting
@@ 1 Introduction
In this work, we tackle the challenging problem of multi-step uncertainty quantification for time series prediction. Our goal is to construct joint prediction regions (JPRs), a generalisation of prediction intervals to sequences of future values. The naive approach of taking the Cartesian product of marginal prediction intervals, each with the desired coverage level $(1-\epsilon)$, does not guarantee the desired global coverage. If future time steps were independent, this approach would result in coverage $(1-\epsilon)^{H}$, where $H$ is the length of the horizon. Although Bonferroni correction Dunn (1961) can adjust for this, it becomes overly conservative as $H$ increases, especially in the presence of temporal dependencies, which, of course, is to be expected in the case of time series data.
@@
Conformal Prediction (CP) (Angelopoulos and Bates, 2022; Vovk et al., 2023; Fontana et al., 2023) offers a distribution-free frequentist methodology for uncertainty quantification with finite-sample coverage guarantees. Specifically, CP ensures that $\Pr(Y\in\mathcal{C}(X^{*}))\geq 1-\epsilon$, where $\mathcal{C}(X^{*})$ is the prediction set (or interval in the univariate case) for a new input $X^{*}$. The theoretical guarantees and model-agnostic nature of CP have spurred its application in diverse areas, including large language models (Quach et al., 2023) and online model aggregation (Gasparin and Ramdas, 2024). CP is readily applicable to any algorithm that provides a (non-)conformity score, a measure of how well a data point aligns with the rest of the dataset. The key assumption is exchangeability, which is satisfied for independent and identically distributed (IID) data. Extensions of CP exist for dependent data settings (Chernozhukov et al., 2018; Barber et al., 2023; Xu and Xie, 2023), but primarily focus on single-step predictions, and the extension to multistep prediction is not trivial.
@@
CP offers finite-sample guarantees, making it an attractive approach to building joint prediction regions (JPRs). However, the exchangeability assumption of CP is always violated in time series data. Moreover, the multivariate residuals ($H$ in total —- one residual per time step) in a multi-step ahead prediction pose a challenge, as for most regression cases, CP relies on the exchangeability of the scalar non-conformity scores and the quantile inversion for calibration. This poses the question of mapping multivariate residuals to a scalar value. Whilst single-step prediction allows for simple residual sorting and the necessary quantile inversion, this approach fails for multi-step scenarios due to the multidimensionality of the residuals. We address the exchangeability issue in the inductive conformal setting by showing an extension of the work of Chernozhukov et al. (2018) from the transductive (full) to the inductive (split) CP setting. Full CP requires fitting many models, whilst inductive CP only requires fitting a single model. The approach of Chernozhukov et al. (2018) leverages specific index permutations of a single time series, providing a distribution of residuals for each permuted time series. We further introduce a novel non-conformity score designed to map multi-dimensional residuals to univariate quantities, enabling the construction of valid JPRs. The resulting framework, JANET (Joint Adaptive predictioN-region Estimation for Time-series), allows for inductive conformal prediction of JPRs in both multiple and single time series settings. When applied to multiple independent time series, JANET guarantees validity as the permutation across independent time series is exchangeable, whilst, for single time series, it provides approximate validity under weak assumptions on the non-conformity scores as long as transformations (permutations within a single series) of the data are a meaningful approximation for a stationary series.
@@
Our key contributions are: (i) formally generalising the framework in Chernozhukov et al. (2018) to the inductive setting for computational efficiency whilst maintaining approximate coverage guarantees; (ii) design of non-conformity measures that effectively control the $K$-familywise error (a generalisation of the familywise error) Lehmann and Romano (2005); Wolf and Wunderli (2013), whilst accounting for time horizon and historical context; and (iii) empirical demonstration of JANET’s (finite-sample) coverage guarantees, computational efficiency, and adaptability to diverse time series scenarios.
2 Related work
Relaxing exchangeability
Recent research has shown a growing interest in adapting conformal prediction (CP) to non-exchangeable data. Early work by Vovk et al. (2005) explored relaxing the exchangeability assumption using Mondrian CP, which divides observations into exchangeable groups. Dunn et al. (2022) built upon this idea to share strengths between groups in hierarchical modelling. Tibshirani et al. (2020) and (Podkopaev and Ramdas, 2021) addressed covariate and label shifts, respectively, by reweighting data based on likelihood ratios. Similarly, Lei and Candès (2021) and Candès et al. (2023) applied reweighting for predictive inference in causal inference and survival analysis, while (Fannjiang et al., 2022) focused on controlling covariate shift by statisticians. However, these methods address heterogeneity rather than the serial dependence found in time series.
One-step or multivariate prediction
Gibbs and Candès (2021) tackled distribution shifts in an online manner by adapting coverage levels based on comparisons of current coverage with desired coverage. Zaffran et al. (2022) extended this work with online expert aggregation. Gibbs and Candès (2023) later introduced an expert selection scheme to guide update step sizes. These works typically require a gradient-based approach to learn a model that adapts to the coverage. Barber et al. (2023) generalised pure conformal inference for dependent data, using fixed weights for recent training examples to account for distributional drift. Xu and Xie (2023) employed predictor ensembling, assuming exchangeability but with asymptotic guarantees. Chernozhukov et al. (2018) leveraged randomisation inference (Romano, 1990) to generalise full conformal inference to serial dependence, achieving valid coverage under exchangeability and approximate validity with serial dependence. Notably, these methods primarily focus on single-step predictions for univariate series. A similar extension to the Inductive setting is provided in Diquigiovanni et al. (2024) for the functional time series setting. They randomly selected the time indices to form a calibration series, however, such an approach does not preserve the statistical properties of the time series. Additionally, the work is primarily based on applying to the time series and lacks formalisation as the generalised inductive conformal predictors. In contrast, we split the sequence into two such that we preserve the statistical properties of the time series; our method formally extends Chernozhukov et al. (2018) to inductive conformal prediction and multi-step scenarios, including multivariate time series.
Other notable works, Diquigiovanni et al. (2021); Ajroldi et al. (2023); Xu et al. (2024) constructed prediction bands for multivariate functional data, functional surfaces and ellipsoidal regions for multivariate time series, respectively, but only for single-step predictions.
Multi-step prediction
Alaa and van der Schaar (2020) applied the jackknife method to RNN-based networks for multi-step prediction, with theoretical coverage of $1-2\epsilon$ at significance level $\epsilon$. Stankeviciute et al. (2021) assumed conditional IID prediction errors in multi-output models, using Bonferroni correction (Haynes, 2013) to achieve desired coverage. However, their approach can be overly conservative with increasing prediction horizons and may not hold when model assumptions are violated. Additionally, they require multi-output prediction models. Cleaveland et al. (2024) utilised linear complementary programming and an additional dataset to optimise the parameters of a multi-step prediction error model. In recent work, Sun and Yu (2024), building on Messoudi et al. (2021), employed copulas to adjust for temporal dependencies, enabling multi-step and autoregressive predictions for multivariate time series filling in the deficit from other works. However, their method requires two calibration sets and gradient-based optimisation, making it data-inefficient. Furthermore, unlike our proposed framework JANET, their approach requires multiple independent time series (just like the preceding works) and cannot adapt prediction regions based on historical context.
A concurrent work Zhou et al. (2024) primarily focussed on single-step prediction adapted Gibbs and Candès (2021) to account for heterogeneous trajectories in multiple time series settings. However, they require multiple calibration sets and do not show how to control $K-$familywise error when dealing with multi-step prediction.
Non-conformal prediction regions
Beyond conformal methods, bootstrapping provides an alternative method for constructing joint prediction regions especially when one only has a single time series. Staszewska-Bystrova (2013) generates $B$ bootstraps and finds a heuristic-based prediction region from the bootstrapped predictive sequences, however, the method only provides have asymptotic guarantees. Wolf and Wunderli (2013) creates JPRs that are asymptotically valid and can control the $K$-familywise error. However, given that this method is based on bootstrapping, it has a large computational cost that is infeasible when working with neural networks. Our work can be seen as conformalisation of Wolf and Wunderli (2013); English (2024) without relying on bootstrapping and with adaptive prediction regions. Additionally, as a conformal method, JANET can be readily applied to time series classification tasks which is not apparent for bootstrap-based methods.
@@ 3 Background on conformal prediction
In this section, we provide an overview of conformal prediction for IID data in the context of regression tasks. In the next section, we specify the CP for the time-series setting.
Full conformal predictors assess the non-conformity of a test sample $x^{*}$ and a postulated target $y$ to a training set by running the underlying algorithm $(n+1)c$ times, where $n$ is the number of training samples and $c$ is the number of points along a discretised grid of the target space ($\mathcal{Y}$). This is done to construct prediction sets by inversion: the set of grid points that best conform according to a desired significance level $\epsilon$. Full CP is prohibitive for compute-intensive underlying training algorithms as it requires refitting a model for each postulated label/grid point. Inductive conformal predictors (ICPs) offer an elegant solution to this problem by training the model only once. We focus on regression tasks with ICPs.
@@
In standard regression, a natural non-conformity score is the absolute residual $a_{i}=|y_{i}-\hat{f}(x_{i})|$. By construction, for any $i$, $y_{i}\in[\hat{f}(x_{i})\pm a_{i}]$; $a_{i}$ is half the interval width that ensures coverage for any new sample (assuming symmetric intervals). In ICP we compute a non-conformity score for each of the $n_{cal}$ samples in the calibration set and then srot them from largest to smallest. Let $a_{(1)}$ denote the largest and $a_{(n_{cal})}$ denote the smallest. Then intervals of the form $(\hat{f}(x_{i})\pm a_{(1)})$ will cover all but one sample from the calibration set and intervals of the form $(\hat{f}(x_{i})\pm a_{(n_{cal})})$ will cover none of the samples from the calibration set (assuming no ties).
Extending this line of thinking, a prediction interval with $(1-\epsilon)$ coverage can be obtained by inverting the quantile of the distribution of non-conformity scores $a_{i}$. To do so we find the $\lfloor\epsilon(n_{cal}+1)\rfloor^{th}$ largest non-conformity score, $q^{a}_{1-\epsilon}:=a_{(\lfloor\epsilon(n_{cal}+1\rfloor)}$ and set this to half of our interval width. For an unseen $Z^{*}=(X^{*},Y^{*})$, we will predict
@@
A drawback of this method is that the intervals are of constant width, $2q^{a}_{1-\epsilon}$. Lei et al. (2018) suggest an alternative non-conformity score that takes into account local information, $r_{i}=\frac{|y_{i}-\hat{f}(x_{i})|}{\hat{s}(x_{i})}$. Here, $\hat{s}(\cdot)$ is also fit on the train set and predicts the conditional mean absolute deviation. Now we can provide locally adaptive prediction intervals for a test sample $Z^{*}=(X^{*},Y^{*})$ with $(1-\epsilon)$ coverage:
@@ 4 Generalised inductive conformal predictors
In this section, we formally describe a generalised framework for inductive conformal predictors based on Generalised CP (Chernozhukov et al., 2018) and randomisation inference (Romano, 1990) applied to time series forecasting. With minor notational adjustments, our generalised ICP extends to multi-step and multivariate time series prediction. In the following sections, we demonstrate how to obtain exact validity guarantees for independent time series (or IID data) and approximate validity guarantees for a single time series. Our task is to forecast $H$ steps into the future conditioned on $T$ steps of history.
@@ 4.1 Multiple time series
Assume we have $n$ independent time series $\mathbf{Z}=\{Z_{k}\}_{k=1}^{n}$ where each individual time series, $Z_{k}$, is an independent realisation from an underlying distribution $\mathbb{P}$ (and within each $Z_{k}$ there is temporal dependence). As usual in ICP, we split $\mathbf{Z}$ into a proper training sequence $\mathbf{Z}_{tr}=\{Z_{k}\}_{k=1}^{n_{tr}}$ and a calibration sequence $\mathbf{Z}_{cal}=\{Z_{n_{tr}+i}\}_{i=1}^{n_{cal}}$. Without loss of generality, we assume the time series are length $T+H$. For each time series, we can define $Z_{k}=\{z_{k,1},z_{k,2},\dots,z_{k,T+H}\}$ as $(X_{k},Y_{k})$, each with $X_{k}=\{z_{k,1},\dots z_{k,T}\}=\{x_{k,1},\dots x_{k,T}\}$ denoting the relevant series’ history, $Y_{k}=\{z_{k,T+1},\dots,z_{k,T+H}\}=\{y_{k,1},\dots,y_{k,H}\}$ being the values at the next $H$ time steps and each $z_{k,j}\in\mathbb{R}^{p}$ (where $p=1$ corresponds to univariate time series and $p>1$ corresponds to multivariate time series). In other words, each time series can be split into the history we use to predict ($X_{k}$) and the target ($Y_{k}$). As in any other ICP setting, we can train a model $\hat{f}:\mathbb{R}^{p\times T}\to\mathbb{R}^{p\times H}$ that predicts $H$ steps into the future based on $T$ steps of history. Then, we can compute non-conformity scores for a non-conformity scoring function $\mathcal{A}$. Note that we can form a distribution over non-conformity scores with one non-conformity score per time series. From there we can invert the quantiles and produce a prediction interval.
@@ 4.2 Single time series
Unlike the case of multiple time series, we only have a single time series and we do not have access to an entire distribution of non-conformity scores—we only have a single score. To address this problem we apply a permutation scheme from Chernozhukov et al. (2018) on the calibration series that provides a distribution over the conformity scores, this is equivalent to the view of randomisation inference employed in (Chernozhukov et al., 2018). We compute the $p$-values as in Definition 1. Now we can treat this collection of permuted time series in the same way we did for multiple time series. Each permutation will provide a non-conformity score and this is how we approximate a distribution of non-conformity scores.
We now assume we have a single time series $Z=\{z_{1},\dots z_{L}\}$ where $L=L_{tr}+L_{cal}$ is the length of the entire single series, $L_{tr},L_{cal}\geq T+H$ and all $z_{i}\in\mathbb{R}^{p}$. We split $Z$ into a train subseries, $Z_{tr}=\{z_{1},\dots,z_{L_{tr}}\}$, and a calibration subseries, $Z_{cal}=\{z_{L_{tr}+1},\dots,z_{L_{tr}+L_{cal}}\}$. $Z_{tr}$ is used to fit our point prediction model $\hat{f}$ and $Z_{cal}$ will be used to calibrate our prediction intervals.
Let $\Pi$ be a set of permutations of the indices $\{1,\dots,L_{cal}\}$. For a permutation $\pi\in\Pi$, let $Z^{\pi}_{cal}=\{z_{L_{tr}\pi(i)}\}_{i=1}^{L_{cal}}$ denote a permuted version of the calibration sequence, $Z_{cal}$. We call $\mathbf{Z}^{\Pi}_{cal}=\{Z^{\pi}_{cal}\}_{\pi\in\Pi}$ the set of permuted version of $Z_{cal}$ under the permutations in $\Pi$. We can partition each permutation of $Z_{cal}$ into $X$ and $Y$ components as in the multiple time series setup where $X$ is the lagged version of $Y$.
@@
Chernozhukov et al. (2018) introduced a permutation scheme to address data dependence. We adapt this scheme and their notation to our setting on $Z_{cal}$. Given a time calibration sequence of length $L_{cal}$, divisible (for simplicity) by a block size $b$, we divide the calibration series into $d=L_{cal}/b$ non-overlapping blocks of $b$ observations each. The $j^{th}$ non-overlapping block (NOB) permutation is defined by the permutation $\Pi_{j,NOB}:\{1,\dots,L_{cal}\}\to\{1,\dots,L_{cal}\}$ where
@@
Figure 1: Visualisation of NOB permutations applied to a single time series $Z$ of length $L=L_{tr}+L_{cal}=10+6=16$ ($z_{1},\dots,z_{10}$ are used for training and $z_{11},\dots,z_{16}$ are reserved for calibration). The first row shows the calibration portion of a time series, $Z$, under the identity permutation, $\pi_{1}$. Subsequent rows show how different permutations rearrange $Z$ for block size, $b=1$. Arrows denote how the objects are permuted. Each permutation in this group rotates the front block to the end.

### v2

@@ Abstract
keywords:
Conformal Prediction, Joint Prediction Region, Time Series Uncertainty Quantification
@@ 1 Introduction
In this work, we tackle the challenging problem of multi-step uncertainty quantification for time series prediction. Our goal is to construct joint prediction regions (JPRs), a generalisation of prediction intervals to sequences of future values. The naive approach of taking the Cartesian product of marginal prediction intervals, each with the desired coverage level $(1-\epsilon)$, does not guarantee the desired global coverage. If future time steps were independent, this approach would result in coverage $(1-\epsilon)^{H}$, where $H$ is the length of the horizon. Although techniques such as Bonferroni correction [14] can adjust for this, it becomes overly conservative as $H$ increases, especially in the presence of temporal dependencies, which, of course, is to be expected in the case of time series data.
Existing approaches to uncertainty quantification in time series, such as bootstrap-based prediction intervals or Bayesian forecasting techniques (e.g., dropout [19] or variational inference), each come with key limitations. Bootstrap methods offer only asymptotic coverage guarantees and often break down under small-sample or dependent data settings. Bayesian methods, on the other hand, require the specification of priors and frequently produce miscalibrated intervals when the model is misspecified.
@@
Unlike these approaches, Conformal Prediction (CP) [4, 41, 34] offers a distribution-free frequentist methodology for uncertainty quantification with finite-sample coverage guarantees. Specifically, CP ensures that $\Pr(Y\in\mathcal{C}(X^{*}))\geq 1-\epsilon$, where $\mathcal{C}(X^{*})$ is the prediction set (or interval in the univariate case) for a new input $X^{*}$. The theoretical guarantees and model-agnostic nature of CP have spurred its application in diverse areas, including large language models [31] and online model aggregation [20]. CP is readily applicable to any algorithm that provides a (non-)conformity score, a measure of how well a data point aligns with the rest of the dataset. The key assumption is exchangeability, which is satisfied for independent and identically distributed (IID) data. Extensions of CP exist for dependent data settings [10, 7, 43], but primarily focus on single-step predictions, and the extension to multistep prediction is not trivial.
@@
CP offers finite-sample guarantees, making it an attractive approach to building joint prediction regions (JPRs). However, the exchangeability assumption of CP is always violated in time series data. Moreover, the multivariate residuals ($H$ in total — one residual per time step) in a multi-step ahead prediction pose a challenge, as for most regression cases, CP relies on the exchangeability of the scalar non-conformity scores and the quantile inversion for calibration. This poses the question of mapping multivariate residuals to a scalar value. Whilst single-step prediction allows for simple residual sorting and the necessary quantile inversion, this approach fails for multi-step scenarios due to the multidimensionality of the residuals. We address the exchangeability issue in the inductive conformal setting by showing an extension of the work of [10] from the transductive (full) to the inductive (split) CP setting. Full CP requires fitting many models, whilst inductive CP only requires fitting a single model. The approach of [10] leverages specific index permutations of a single time series, providing a distribution of residuals for each permuted time series. We further introduce a novel non-conformity score designed to map multi-dimensional residuals to univariate quantities, enabling the construction of valid JPRs. The resulting framework, JANET (Joint Adaptive predictioN-region Estimation for Time-series), allows for inductive conformal prediction of JPRs in both multiple and single time series settings. When applied to multiple independent time series, JANET guarantees validity as the permutation across independent time series is exchangeable, whilst, for single time series, it provides approximate validity under weak assumptions on the non-conformity scores as long as transformations (permutations within a single series) of the data are a meaningful approximation for a stationary series.
@@
Our key contributions are: (i) formally generalising the framework in [10] to the inductive setting for computational efficiency whilst maintaining approximate coverage guarantees; (ii) design of non-conformity measures that effectively control the $K$-familywise error (a generalisation of the familywise error) [26, 42], whilst accounting for time horizon and historical context; and (iii) empirical demonstration of JANET’s (finite-sample) coverage guarantees, computational efficiency, and adaptability to diverse time series scenarios.
@@ 2 Background on conformal prediction
In this section, we provide an overview of conformal prediction for IID data in the context of regression tasks. In the next section, we specify CP for the time series setting.
Full conformal prediction constructs prediction sets by evaluating how well a hypothesised label $y$ for a new input $X^{*}$ conforms with the training data. For each candidate value $y$ from a discretised grid over the target space $\mathcal{Y}$, the method temporarily augments the training set with the test pair $(X^{*},y)$, forming a dataset of size $n+1$. A model is then retrained on this augmented dataset to compute a non-conformity score for each of the $n+1$ points, including the test point. This process is repeated across all $c$ grid points, resulting in $(n+1)\times c$ model fits. For each $y$, a p-value is computed by ranking the test point’s score among those of the training samples under the exchangeability assumption. The prediction set includes all $y\in\mathcal{Y}$ such that the corresponding p-value exceeds the significance level $\epsilon$; this inversion of the conformity scores yields a valid $(1-\epsilon)$ prediction set. Full CP is prohibitive for compute-intensive underlying training algorithms as it requires $(n+1)\times c$ model fits. Inductive conformal predictors (ICPs), explained below, offer an elegant solution to this problem by training the model only once. We focus on regression tasks with ICPs.
@@
In standard regression, a natural non-conformity score is the absolute residual $a_{i}=|y_{i}-\hat{f}(x_{i})|$. By construction, for any $i$, $y_{i}\in[\hat{f}(x_{i})\pm a_{i}]$; $a_{i}$ is half the interval width that ensures coverage for any new sample (assuming symmetric intervals). In ICP we compute a non-conformity score for each of the $n_{cal}$ samples in the calibration set and then sort them from largest to smallest. Let $a_{(1)}$ denote the largest and $a_{(n_{cal})}$ denote the smallest. Then intervals of the form $(\hat{f}(x_{i})\pm a_{(1)})$ will cover all but one sample from the calibration set and intervals of the form $(\hat{f}(x_{i})\pm a_{(n_{cal})})$ will cover none of the samples from the calibration set (assuming no ties).
Extending this line of thinking, a prediction interval with $(1-\epsilon)$ coverage can be obtained by inverting the quantile of the distribution of non-conformity scores $a_{i}$. To do so we find the $\lfloor\epsilon(n_{cal}+1)\rfloor^{th}$ largest non-conformity score, $q^{a}_{1-\epsilon}:=a_{(\lfloor\epsilon(n_{cal}+1\rfloor)}$ and set this to half of our interval width. For an unseen $Z^{*}=(X^{*},Y^{*})$, we will provide a prediction interval of the form
@@
A drawback of this method is that the intervals are of constant width, $2q^{a}_{1-\epsilon}$. [27] suggest an alternative non-conformity score that takes into account local information, $r_{i}=\frac{|y_{i}-\hat{f}(x_{i})|}{\hat{s}(x_{i})}$. Here, $\hat{s}(\cdot)$ is also fit on the train set and predicts the conditional mean absolute deviation. Now we can provide locally adaptive prediction intervals for a test sample $Z^{*}=(X^{*},Y^{*})$ with $(1-\epsilon)$ coverage:
@@ 3 Generalised inductive conformal predictors
In this section, we formally describe a generalised framework for inductive conformal predictors based on Generalised CP [10] and randomisation inference [33] applied to time series forecasting. With minor notational adjustments, our generalised ICP extends to multi-step and multivariate time series prediction. In the following sections, we demonstrate how to obtain exact validity guarantees for independent time series (or IID data) and approximate validity guarantees for a single time series. Our task is to forecast $H$ steps into the future conditioned on $T$ steps of history.
@@ 3.1 Multiple time series
Assume we have $n$ independent time series $\mathbf{Z}=\{Z_{k}\}_{k=1}^{n}$ where each individual time series, $Z_{k}$, is an independent realisation from an underlying distribution $\mathbb{P}$ (and within each $Z_{k}$ there is temporal dependence). As usual in ICP, we split $\mathbf{Z}$ into a proper training sequence $\mathbf{Z}_{tr}=\{Z_{k}\}_{k=1}^{n_{tr}}$ and a calibration sequence $\mathbf{Z}_{cal}=\{Z_{n_{tr}+i}\}_{i=1}^{n_{cal}}$. For notational convenience, we assume the time series are of length $T+H$. For each time series, we can define $Z_{k}=\{z_{k,1},z_{k,2},\dots,z_{k,T+H}\}$ as $(X_{k},Y_{k})$, each with $X_{k}=\{z_{k,1},\dots z_{k,T}\}=\{x_{k,1},\dots x_{k,T}\}$ denoting the relevant series’ history, $Y_{k}=\{z_{k,T+1},\dots,z_{k,T+H}\}=\{y_{k,1},\dots,y_{k,H}\}$ being the values at the next $H$ time steps and each $z_{k,j}\in\mathbb{R}^{p}$ (where $p=1$ corresponds to univariate time series and $p>1$ corresponds to multivariate time series). In other words, each time series can be split into the history we use to predict ($X_{k}$) and the target ($Y_{k}$). As in any other ICP setting, we can train a model $\hat{f}:\mathbb{R}^{p\times T}\to\mathbb{R}^{p\times H}$ that predicts $H$ steps into the future based on $T$ steps of history. Then, we can compute non-conformity scores for a non-conformity scoring function $\mathcal{A}$. Note that we can form a distribution over non-conformity scores with one non-conformity score per time series. From there we can invert the quantiles and produce a prediction interval.
@@ 3.2 Single time series
Unlike the case of multiple time series, we only have a single time series and we do not have access to an entire distribution of non-conformity scores—we only have a single score. To address this problem we apply a permutation scheme from [10] on the calibration series that provides a distribution over the conformity scores, this is equivalent to the view of randomisation inference employed in [10].
@@
[10] introduced a permutation scheme to address data dependence. We adapt this scheme and their notation to our setting on $Z_{cal}$. Given a time calibration sequence of length $L_{cal}$, divisible (for simplicity) by a block size $b$, we divide the calibration series into $d=L_{cal}/b$ non-overlapping blocks of $b$ observations each. The $j^{th}$ non-overlapping block (NOB) permutation is defined by the permutation $\Pi_{j,NOB}:\{1,\dots,L_{cal}\}\to\{1,\dots,L_{cal}\}$ where
@@
Now we can treat this collection of permuted time series in the same way we did for multiple time series. Each permutation will provide a non-conformity score and this is how we approximate a distribution of non-conformity scores.
Formally, we now assume we have a single time series $Z=\{z_{1},\dots z_{L}\}$ where $L=L_{tr}+L_{cal}$ is the length of the entire single series, $L_{tr},L_{cal}\geq T+H$ and all $z_{i}\in\mathbb{R}^{p}$. We split $Z$ into a train subseries, $Z_{tr}=\{z_{1},\dots,z_{L_{tr}}\}$, and a calibration subseries, $Z_{cal}=\{z_{L_{tr}+1},\dots,z_{L_{tr}+L_{cal}}\}$. $Z_{tr}$ is used to fit our point prediction model $\hat{f}$ and $Z_{cal}$ will be used to calibrate our prediction intervals.
Let $\Pi$ be a set of permutations of the indices $\{1,\dots,L_{cal}\}$. For a permutation $\pi\in\Pi$, let $Z^{\pi}_{cal}=\{z_{L_{tr}\pi(i)}\}_{i=1}^{L_{cal}}$ denote a permuted version of the calibration sequence, $Z_{cal}$. We call $\mathbf{Z}^{\Pi}_{cal}=\{Z^{\pi}_{cal}\}_{\pi\in\Pi}$ the set of permuted version of $Z_{cal}$ under the permutations in $\Pi$. We can partition each permutation of $Z_{cal}$ into $X$ and $Y$ components as in the multiple time series setup where $X$ is the lagged version of $Y$. We compute the $p$-values as in Definition 1.
Definition 1 (Randomised $p$-value).
We define the randomised $p$-value for the postulated label $y$ as:
$\hat{p}=\hat{p}(y):=\frac{1}{d}\sum_{j=1}^{d}\mathbf{1}\left(\mathcal{A}\left(Z_{tr},Z_{cal}^{\pi_{j}},y\right)\geq\mathcal{A}\left(Z_{tr},Z_{cal},y\right)\right)$
@@
Figure 1: Illustration of Non-Overlapping Block (NOB) permutations applied to the calibration portion of a single time series $Z$ of length $L=L_{tr}+L_{cal}=16$. The first 10 points ($z_{1},\dots,z_{10}$) are used for training, and the final 6 points ($z_{11},\dots,z_{16}$) form the calibration segment. Each row in the figure corresponds to a different permuted version of the calibration sequence $Z_{cal}$, where a block size $b=1$ is used. The top row represents the identity permutation (i.e., no rotation). Each subsequent row is obtained by rotating the calibration sequence one block to the left compared to the row above it. The arrows indicate how elements are rearranged to produce the next new permutation. For example, the second row shows the permutation $\pi_{2}$ applied to the identity (first row), and so on.